1. Using the mass of m-chloropropiophenone found in the procedure (0.2g), calculate the moles of m-chloropropiophenone used. Show your work.

2. A solution of bromine in dichloromethane contains 5.0% bromine by volume. The density of pure, liquid bromine is 3.12 g/mL. Using this information and the moles of m-chloropropiophenone determined above, calculate the volume of the bromine solution required to react with the m-chloropropiophenone. Show your work.
Answer #1

1.

Mass of m-chloropropiophenone used = 0.2 g

Molar mass of m-chloropropiophenone = 168.62 g/mol

Moles of m-chloropropiophenone = 0.2/168.62 = 0.0012 mol

2.

1 mol of m-chloropropiophenone reacts with 1 mol of bromine to give 1 mol of the product.

Therefore, 0.0012 mol of m-chloropropiophenone reacts with 0.0012 mol of bromine to give 0.0012 mol of the product.

Molar mass of Br2 = 160 g/mol

Mass of neat Br2 required for the reaction = molar mass of Br2 x moles of Br2 required

                                                                   = 160 x 0.0012

                                                                   = 0.192 g

Density of Br2 = 3.12 g/mL

Thus, the volume of neat Br2 required for the reaction = mass of Br2 required for the reaction/ density of Br2

                                                                                                      = 0.192/3.12

                                                                                     = 0.062 mL

Now, let us say the volume of the bromine solution required to react with the m-chloropropiophenone = V mL

Hence,

V x 5/100 = 0.062

or, V = (0.062 x 100)/5

        = 1.24 mL

Therefore, the volume of the bromine solution required to react with the m-chloropropiophenone = 1.24 mL